# OneBit: Towards Extremely Low-bit Large Language Models

### v2

@@ Abstract
Model quantification uses low bit-width values to represent the weight matrices of models, which is a promising approach to reduce both storage and computational overheads of deploying highly anticipated LLMs. However, existing quantization methods suffer severe performance degradation when the bit-width is extremely reduced, and thus focus on utilizing 4-bit or 8-bit values to quantize models. This paper boldly quantizes the weight matrices of LLMs to 1-bit, paving the way for the extremely low bit-width deployment of LLMs. For this target, we introduce a 1-bit quantization-aware training (QAT) framework named OneBit, including a novel 1-bit parameter representation method to better quantize LLMs as well as an effective parameter initialization method based on matrix decomposition to improve the convergence speed of the QAT framework. Sufficient experimental results indicate that OneBit achieves good performance (at least 83% of the non-quantized performance) with robust training processes when only using 1-bit weight matrices. Code and checkpoints are available: https://github.com/xuyuzhuang11/OneBit
OneBit: Towards Extremely Low-bit Large Language Models
Yuzhuang Xu1, Xu Han1, Zonghan Yang1, Shuo Wang1 Qingfu Zhu2, Zhiyuan Liu1, Weidong Liu1, Wanxiang Che2,thanks: Corresponding author 1Department of Computer Science & Technology, Tsinghua University, Beijing, China 2Research Center for Social Computing and Information Retrieval, Harbin Institute of Technology, Harbin, China xyz21thu@gmail.com, car@ir.hit.edu.cn
@@ 1 Introduction
Transformer (Vaswani et al., 2017) has emerged as the pivotal architecture in large language models (LLMs), fundamentally reshaping the approach to natural language processing in deep learning era (Bubeck et al., 2023; Touvron et al., 2023; Bisk et al., 2020). Despite their popularity, deploying transformer-based LLMs presents significant challenges due to their computational intensity and considerable memory requirements as the parameters of LLMs become more and more. For instance, even moderately-sized LLMs like LLaMA-13B (Touvron et al., 2023) require around 26GB of memory to load its all parameters in FP16 format. Such overheads make deploying LLMs difficult beyond mid-to-high-end GPUs like the A100, let alone on mobile devices. The high demand for resources not only drives up usage costs, but also restricts their wider application.
@@
Numerous efforts (Dettmers et al., 2022; Frantar et al., 2022; Frantar and Alistarh, 2023) have been devoted to reducing the computational and memory overheads of LLMs, while still preserving most of their original model capabilities. Among these efforts, quantization has gained widespread attention, particularly Post-Training Quantization (PTQ), benefitted from its lower transferring costs. Seminal studies such as GPTQ (Frantar et al., 2022), SpQR (Dettmers et al., 2023b), and AWQ (Lin et al., 2023) successfully compress the weight matrices of LLMs to 4-bit values while maintaining the main abilities of LLMs. Efficient quantization represents significant advances in LLM optimization, by achieving a balance between time and space efficiency as well as model performance.
Figure 1: The perplexity (lower scores mean better performance) of existing widely-used low-bit quantization methods on LLaMA-7B, reported on Wikitext2 (Merity et al., 2016). All the examined previous approaches suffer from significant performance degradation when quantizing models to 2-bit values. Our 1-bit quantization method can outperform these 2-bit baselines.
@@
Unfortunately, the efficacy of PTQ rapidly diminishes when the quantization bit-width is extremely low, as shown in Figure 1. Existing PTQ methods managed to compress weight matrices down to at least 3-bit (Dettmers and Zettlemoyer, 2023). Recent researches hope to leverage Quantization-Aware Training (QAT) to overcome the bottlenecks faced by PTQ. LLM-QAT (Liu et al., 2023) introduces a few learnable parameters into the quantization process, achieving notable results. OmniQuant (Shao et al., 2023), integrating learnable equivalent transformation, presents promising results in 2-bit quantization. However, existing methods decline when compressing model weights to 1 bit, struggling to maintain effectiveness. This mainly stems from the drastic precision loss at extremely low bit-width in weight matrix $\mathbf{W}$, significantly increasing loss in linear projection $\mathbf{WX}$, which is the core operator within LLMs.
@@
In this paper, we propose a novel Linear layer and Sign-Value-Independent Decomposition (SVID) for weight matrices to represent LLMs using approximately 1-bit values. In SVID, each original high-bit weight matrix is decomposed into one sign matrix (±$\pm 1$) and two value vectors. The value vectors provide necessary floating-point precision in linear projection at little cost and help the model to be trained easily. The sign matrix maintains the high rank of the original weight matrix with a small space cost, thereby preserving high information capacity. SVID offers a better parameter initialization for 1-bit models and we employ quantization-aware knowledge distillation to transfer the capabilities of the original model to the proposed 1-bit counterpart. Experiments demonstrate that our method performs well at the W1A16 (1-bit weight and 16-bit activation) quantization level. Furthermore, our 1-bit model is more amenable to training and knowledge transfer than previous works. In summary, the contributions of this work are 3-fold:
@@ 2.1 Large Language Model Compression
Quantization, pruning, and knowledge distillation (KD) are the mainstream methods for model compression. Quantization compresses model weights into low-bit values (Frantar et al., 2022; Lin et al., 2023; Dettmers et al., 2023a). For data type alignment in computation and reducing memory, it also involves quantizing activation (Dettmers et al., 2022; Xiao et al., 2023) and key-value cache (Shao et al., 2023). Pruning simplifies model complexity by removing unimportant weights or modules, thereby sparsifying the original larger models (Frantar and Alistarh, 2023; Sun et al., 2023; Ma et al., 2023). KD trains a smaller student model under the guidance of a larger teacher model (Hsieh et al., 2023; Agarwal et al., 2023; Hsieh et al., 2023), achieving the purpose of compressing the larger one. Beyond these methods, low-rank factorization approximates the original weight matrix $\mathbf{W}$ with the product of two lower-rank matrices (Xu et al., 2023) and also achieves promising results. Our work belongs to quantization, using KD for knowledge transfer from the original LLM and uniquely focusing on extremely low bit-width quantization. More details about model compression can refer to existing survies  (Wan et al., 2023; Zhu et al., 2023).
@@ 2.2 Large Language Model Quantization
PTQ directly converts trained models into lower-bit counterparts using accurate solvers and limited calibration data without additional training. Typically, GPTQ (Frantar et al., 2022) row-wisely quantizes weight matrices and adjusts remaining weights to compensate for the precision loss caused by quantization, achieving nearly lossless 4-bit weight quantization. Moreover, numerous studies observed the effect of “outliers” in quantization (Dettmers et al., 2022; Kim et al., 2023b; Lin et al., 2023). LLM.int8() (Dettmers et al., 2022) suggests mixed-precision decomposition to ensure the accuracy of a few outliers in activations. SmoothQuant (Xiao et al., 2023) reduces the difficulty of quantization by smoothing the outliers of activation. SpQR (Dettmers et al., 2023b) identifies sensitive weights to ensure their precision, while quantizing other weights to lower bit-width.
QAT integrates quantization steps within the model, applying them during training or fine-tuning. It allows the model to better adapt to the reduced precision induced by quantization, leading to improved performance compared to PTQ. LLM-QAT (Liu et al., 2023) introduces a small number of learnable parameters into quantization and employs KD using data generated by the original model itself. OmniQuant (Shao et al., 2023; we classify it as QAT) further introduces learnable equivalent transformation, achieving acceptable results in 2-bit weight quantization. PEQA (Kim et al., 2023a) and QLoRA (Dettmers et al., 2023a) focus on fine-tuning a limited number of extra parameters to mitigate the precision loss caused by sub-4bit weight quantization. Our work is closely related to QAT, but due to the unique challenges posed by 1-bit quantization, our representation and initialization methods of quantized weights are distinct from any existing work.
@@ 3.1 Background
where $s$ denotes the quantization scale parameter, $z$ denotes the zero point parameter, and $N$ is the quantization bit-width. $\mathrm{Clip}(\cdot)$ truncates the result in the range of 00 to $12^{N}-$ 1. With the bit-width being lower and lower, the quantization grid also becomes sparser. When we quantize a LLM to 1-bit values, there are only 2 available numbers to be chosen in the quantized model. Existing study (Dettmers and Zettlemoyer, 2023) points out that quantization based on the RTN method may get their best performance at the 4-bit level. Further compressing like quantizing models to 2-bit values would result in a substantial degradation (Shao et al., 2023) as shown in Figure 1.
@@
Figure 2: The main idea of our method OneBit. The left is the original FP16 Linear Layer, in which both the activation $\mathbf{X}$ and the weight matrix $\mathbf{W}$ are in FP16 format. The right is our proposed architecture. Only value vectors $\mathbf{g}$ and $\mathbf{h}$ are in FP16 format and the weight matrix consists of ±$\pm 1$ instead.
@@
Due to the severe precision loss of 1-bit weight quantization, converting weight matrices in Linear layers directly from FP32/16 to 1-bit format based on RTN is challenging. Wang et al. (2023) explore this possibility by studying the capabilities of purely 1-bit weight matrices, training the 1-bit model from scratch. In the W1A16 setting, their Linear layers are designed as
@@
where $\mathbf{W}$ denotes the quantized weight matrix with the shape $m\times n$ and $\mathbf{W}_{\pm 1}$ denotes the 1-bit quantized matrix. $\mathbf{X}$ is the input of Linear layer and $\mathbf{Y}$ is the output. $\mathrm{Sign}(\cdot)$, $\mathrm{Mean}(\cdot)$ and $\mathrm{Abs}(\cdot)$ functions return the sign matrix, average and absolute value matrix. Unfortunately, this approach reduces computational demands but also leads to a marked decrease in performance Wang et al. (2023).
Inspired by Wang et al. (2023), we also quantize the weight matrix using the function $\mathrm{Sign}(\cdot)$, and the element of the quantized matrix is set to +1 or -1. Moreover, we also notice that although $\mathbf{W}_{\pm 1}$ maintains a high rank of $\mathbf{W}$, the missed floating-point precision still destroys the model performance. Therefore, different from previous work, we introduce two value vectors with an FP16 format to compromise the precision loss in the quantization process. Our proposed Linear layers are designed as
@@
where $\mathbf{g}$ and $\mathbf{h}$ are the two FP16 value vectors. Note that we specify the calculation order using brackets in Eq. (3) for minimizing the time and space cost. The main difference between Wang et al. (2023) and OneBit is the extra parameter $\mathbf{g}$ and $\mathbf{h}$. Even if additional parameters are brought in, the benefits far outweigh its small cost. For instance, when we quantize one weight matrix with the shape $4096\times 4096$, the average bit-width of the quantized result is 1.0073. See A.6 for the details.
@@ 3.3 Sign-Value-Independent Decomposition
We can employ some widely used matrix decomposition methods to perform the rank-1 approximation, such as SVD (Beltrami, 1990) and NMF (Paatero and Tapper, 1994).
@@
We prove this approximation in Appendix A.1. This bridges the gap between the architecture of the quantized model and its original weights. It indicates that if we assign $\mathbf{W}_{\mathrm{sign}}$ to $\mathbf{W}_{\pm 1}$, $\mathbf{a}^{\mathrm{T}}$ to $\mathbf{h}$ and $\mathbf{b}^{\mathrm{T}}$ to $\mathbf{g}$, the quantized model is an approximate initialization of the original model. Moreover, compared to restoring the original matrix $\mathbf{W}$ first (such as in Eq. (4)), the computational order in Eq. (5) saves approximately one matrix $\mathbf{W}$ in FP16 format in memory as there is no need to restore FP16 $\mathbf{W}$.
@@ 3.4 Knowledge Transfer
We employ quantization-aware knowledge distillation to transfer knowledge from the original model (i.e. teacher model) to the quantized one (i.e. student model). In the student model, the element in matrix $\mathbf{W}$ and vectors $\mathbf{g}$/$\mathbf{h}$ in Eq. (3) will be trained. We use cross-entropy based logits and mean-square-error based hidden state of the full-precision teacher model to direct the quantized student model (Sun et al., 2019). Language modeling loss is not used. The cross-entropy is defined as
@@
where $\alpha$ is the hyper-parameter that balances the importance of the cross-entropy loss and the features in the intermediate layers.
@@ 4.1 Settings
For the training data of our quantization-aware knowledge distillation, we follow Liu et al. (2023) to synthesize corpus using next token generation from the original teacher model. It randomizes the first token from vocabulary and generates the next token iteratively until reaching either the <EOS> token or the maximum length. Specially, the top-1 predictions are selected deterministically for the first 3 to 5 tokens, followed by stochastic sampling for the remaining tokens. We utilized LLaMA-7B to generate a total of 132k data entries, each with a maximum length of 2,048.
@@
Every KD experiment learns the training data over 50 epochs, from which 2048-token segments are selected. We employ NMF in scikit-learn  [^1] https://scikit-learn.org/ to decompose the weight matrices in SVID. The quantized student models are optimized by Adam (Kingma and Ba, 2014) with $\beta_{1}=0.9$, $\beta_{2}=0.98$. The learning rate for all experiments is scheduled by cosine strategy. We use NVIDIA A100 GPUs and maintain FP16 precision while training quantized models. For additional details such as learning rate, please refer to Table 1.
Models learning rate $\mathbf{\alpha}$ GPUs
@@
Wiki2 C4 Winogrande Hellaswag PIQA BoolQ ARC-e ARC-c Avg.
@@
To our knowledge, there is no previous work exploring the 1-bit quantization of LLMs from a knowledge transfer perspective. To this end, we relax the quantization bit-width of baselines to 2 bits (W2A16) while maintaining the W1A16 setting in our method. We compare our method with GPTQ (Frantar et al., 2022), LLM-QAT (Liu et al., 2023) and OmniQuant (Shao et al., 2023). To ensure a fair comparison in terms of space usage, baselines do not employ grouped quantization. Additionally, we included the results of vanilla transformers with FP16 precision as a reference. While the recent work BitNet (Wang et al., 2023) also introduced one 1-bit model architecture, it only focused on training models from scratch. We also analyze its capability to transfer knowledge from the original models in Appendix A.5.
@@
Basically, we evaluate quantized models by testing the perplexity on the validation set, specifically on WikiText2 (Merity et al., 2016) and C4 (Raffel et al., 2020). Lower perplexity indicates that the compressed model is better at preserving the output distribution of the original model. Furthermore, accuracies of zero-shot tasks including Winograde (Sakaguchi et al., 2021), HellaSwag (Zellers et al., 2019), PIQA (Bisk et al., 2020), BoolQ (Clark et al., 2019), and ARC (Clark et al., 2018) are also reported. They evaluate if the capabilities of the original model on downstream tasks are retained. We utilize the open-sourced toolkit “LM-Evaluation-Harness” [^2] https://github.com/EleutherAI/lm-evaluation-harness to perform the perplexity test and all zero-shot tasks.

### v3

@@ Abstract
Model quantification uses low bit-width values to represent the weight matrices of existing models to be quantized, which is a promising approach to reduce both storage and computational overheads of deploying highly anticipated LLMs. However, current quantization methods suffer severe performance degradation when the bit-width is extremely reduced, and thus focus on utilizing 4-bit or 8-bit values to quantize models. This paper boldly quantizes the weight matrices of LLMs to 1-bit, paving the way for the extremely low bit-width deployment of LLMs. For this target, we introduce a 1-bit model compressing framework named OneBit, including a novel 1-bit parameter representation method to better quantize LLMs as well as an effective parameter initialization method based on matrix decomposition to improve the convergence speed of the quantization framework. Sufficient experimental results indicate that OneBit achieves good performance (at least 81% of the non-quantized performance on LLaMA models) with robust training processes when only using 1-bit weight matrices. Code and checkpoints are available at https://github.com/xuyuzhuang11/OneBit
@@ 1 Introduction
Transformer [36] has emerged as the pivotal architecture in large language models (LLMs), fundamentally reshaping the approach to natural language processing in deep learning era [6, 34, 4]. Despite their popularity, deploying transformer-based LLMs presents significant challenges due to their computational intensity and considerable memory requirements as the parameters of LLMs become more and more. For instance, even moderately-sized LLMs like LLaMA-13B [34] require around 26GB of memory to load its all parameters in FP16 format. Such overheads make deploying LLMs difficult beyond mid-to-high-end GPUs like the A100, let alone on mobile devices. The high demand for resources not only drives up usage costs, but also restricts their wider application.
@@
Numerous efforts [10, 14, 13] have been devoted to reducing the computational and memory overheads of LLMs, while still preserving most of their original model capabilities. Among these efforts, quantization has gained widespread attention, particularly Post-Training Quantization (PTQ), benefitted from its lower transferring costs. Seminal studies such as GPTQ [14], SpQR [12], and AWQ [20] successfully compress the weight matrices of LLMs to 4-bit values while maintaining the main abilities of LLMs. Efficient quantization represents significant advances in LLM optimization, by achieving a balance between time and space efficiency as well as model performance.
@@
Unfortunately, the efficacy of PTQ rapidly diminishes when the quantization bit-width is extremely low, as shown in Figure 1. Existing PTQ methods managed to compress weight matrices down to at least 3-bit [9]. Recent researches hope to leverage Quantization-Aware Training (QAT) to overcome the bottlenecks faced by PTQ. LLM-QAT [21] introduces a few learnable parameters into
Figure 1: The perplexity (lower scores mean better performance) of existing widely-used low-bit quantization methods on LLaMA-7B, reported on Wikitext2 [23]. All the examined previous approaches suffer from significant performance degradation when quantizing models to 2-bit values. Our 1-bit quantization method can outperform these 2-bit baselines.
the quantization process, achieving notable results. OmniQuant [30], integrating learnable equivalent transformation, presents promising results in 2-bit quantization. However, existing methods decline when compressing model weights to 1 bit, struggling to maintain effectiveness. This mainly stems from the drastic precision loss at extremely low bit-width representation in weight matrix $\mathbf{W}$, significantly increasing loss in linear projection $\mathbf{WX}$, which is the core operator within LLMs.
@@
In this paper, we propose a novel Linear layer and Sign-Value-Independent Decomposition (SVID) for weight matrices to represent LLMs using approximately 1-bit values. In our novel layer architecture, each original high-bit weight matrix is represented as one sign matrix (±$\pm 1$) and two value vectors. The value vectors provide necessary floating-point precision in linear projection at little cost and help the model to be trained easily. The sign matrix maintains the high rank of the original weight matrix with a small space cost, thereby preserving high information capacity. SVID offers a better parameter initialization for 1-bit models from the non-quantized model and we employ quantization-aware knowledge distillation to transfer the capabilities of the original model to the proposed 1-bit counterpart. Experiments demonstrate that our method performs well at the W1A16 (1-bit weight and 16-bit activation) quantization level. Furthermore, our 1-bit model is more amenable to training and knowledge transfer than previous works. In summary, our contributions are 3-fold:
@@ 2.1 Large Language Model Compression
Quantization, pruning, and knowledge distillation (KD) are the mainstream methods for model compression. Quantization compresses model weights into low-bit values [14, 20, 11]. For data type alignment in computation and reducing memory, it also involves quantizing activation [10, 39] and key-value cache [30]. Pruning simplifies model complexity by removing unimportant weights or modules, thereby sparsifying the original larger models [13, 31, 22]. KD trains a smaller student model under the guidance of a larger teacher model [16, 1, 16], achieving the purpose of compressing the larger one. Beyond these methods, low-rank factorization approximates the original weight matrix $\mathbf{W}$ with the product of two lower-rank matrices [40] and also achieves promising results. Our work belongs to quantization, using KD for knowledge transfer from the original LLM and uniquely focusing on extremely low bit-width quantization. More details about model compression can refer to existing survies  [37, 43].
@@ 2.2 Large Language Model Quantization
PTQ directly converts trained models into lower-bit counterparts using accurate solvers and limited calibration data without additional training. Typically, GPTQ [14] row-wisely quantizes weight matrices and adjusts remaining weights to compensate for the precision loss caused by quantization, achieving nearly lossless 4-bit weight quantization. Moreover, numerous studies observed the effect of “outliers” in quantization [10, 18, 20]. LLM.int8() [10] suggests mixed-precision decomposition to ensure the accuracy of a few outliers in activations. SmoothQuant [39] reduces the difficulty of quantization by smoothing the outliers of activation. SpQR [12] identifies sensitive weights to ensure their precision, while quantizing other weights to lower bit-width.
QAT integrates quantization steps within the model, applying them during training or fine-tuning. It allows the model to better adapt to the reduced precision induced by quantization, leading to improved performance compared to PTQ. LLM-QAT [21] introduces a small number of learnable parameters into quantization and employs KD using data generated by the original model itself. OmniQuant (30; we classify it as QAT) further introduces learnable equivalent transformation, achieving acceptable results in 2-bit weight quantization. Contemporary work QuIP# [35] combines randomized Hadamard transform, vector quantization techniques, and fine-tuning to achieve better performance in 2-bit level. PEQA [17] and QLoRA [11] focus on fine-tuning a limited number of extra parameters to mitigate the precision loss caused by sub-4bit weight quantization. Our work is closely related to QAT, but due to the unique challenges posed by 1-bit quantization, our representation and initialization methods of quantized weights are distinct from any existing work.
@@ 3.1 Background
where $s$ denotes the quantization scale parameter, $z$ denotes the zero point parameter, and $N$ is the quantization bit-width. $\mathrm{Clip}(\cdot)$ truncates the result in the range of 00 to $12^{N}-$ 1. With the bit-width being lower and lower, the quantization grid also becomes sparser. When we quantize a LLM to 1-bit values, there are only 2 available numbers to be chosen in the quantized model. Existing study [9] points out that quantization based on the RTN method may get their best performance at the 4-bit level. Further quantizing to 2-bit values following this paradigm would result in a substantial degradation [30] as shown in Figure 1.
@@
Figure 2: The main idea of our method OneBit. The left is the original FP16 Linear Layer, in which both the activation $\mathbf{X}$ and the weight matrix $\mathbf{W}$ are in FP16 format. The right is our proposed architecture. Only value vectors $\mathbf{g}$ and $\mathbf{h}$ are in FP16 format, and the weight matrix consists of ±$\pm 1$ instead, which can be represented in INT1.
@@
Due to the severe precision loss of 1-bit weight quantization, converting weight matrices in Linear layers directly from FP32/16 to 1-bit format based on RTN is challenging. Wang et al. [38] explore this possibility by studying the capabilities of purely 1-bit weight matrices, training the 1-bit model from scratch. In the W1A16 setting, their Linear layers are designed as
@@
where $\mathbf{W}$ denotes the quantized weight matrix with the shape $m\times n$ and $\mathbf{W}_{\pm 1}$ denotes the 1-bit quantized matrix. $\mathbf{X}$ is the input of Linear layer and $\mathbf{Y}$ is the output. $\mathrm{Sign}(\cdot)$, $\mathrm{Mean}(\cdot)$ and $\mathrm{Abs}(\cdot)$ functions return the sign matrix, average and absolute value matrix. Unfortunately, this approach reduces computational demands but also leads to a marked decrease in performance [38]. Moreover, due to training difficulties, experiments show that this method is challenging to use for quantizing existing models and can only be applied to training models from scratch.
Inspired by Wang et al. [38], we also quantize the weight matrix using the function $\mathrm{Sign}(\cdot)$, and the element of the quantized matrix is set to +1 or -1 as well. Moreover, we also notice that although $\mathbf{W}_{\pm 1}$ maintains a high rank of $\mathbf{W}$, the missed floating-point precision still destroys the model performance. Therefore, different from previous work, we introduce 2 value vectors with an FP16 format to compromise the precision loss in the quantization process. During training, our proposed Linear layers are designed as
@@
where $\mathbf{g}$ and $\mathbf{h}$ are the two FP16 value vectors. During inference, $\mathbf{W}_{\pm 1}$ is packed with an INT1 format, and $\mathrm{Sign}(\cdot)$ will not be used, as shown in Figure 2. Note that we specify the calculation order using brackets in Eq. (3) for minimizing the time and space cost. The main difference between Wang et al. [38] and OneBit is the extra parameter $\mathbf{g}$ and $\mathbf{h}$. Even if additional parameters are brought in, the benefits far outweigh its small cost. For instance, when we quantize one weight matrix with the shape $4096\times 4096$, the average bit-width of the quantized result is 1.0073. See A.7 for the details.
@@ 3.3 Sign-Value-Independent Decomposition
We can employ some widely used matrix decomposition methods to perform the rank-1 approximation, such as SVD [2] and NMF [25].
@@
We prove this approximation in Appendix A.1. This bridges the gap between the architecture of the quantized model and its original weights. It indicates that if we assign $\mathbf{W}_{\mathrm{sign}}$ to $\mathbf{W}_{\pm 1}$, $\mathbf{a}^{\mathrm{T}}$ to $\mathbf{h}$ and $\mathbf{b}^{\mathrm{T}}$ to $\mathbf{g}$, the quantized model is an approximate initialization of the original model. Moreover, compared to restoring the original matrix $\mathbf{W}$ first (such as in Eq. (4)), the computational order in Eq. (5) saves approximately one matrix $\mathbf{W}$ in FP16 format in memory as there is no need to restore $\mathbf{W}$ in FP16 format.
@@ 3.4 Knowledge Transfer
We employ quantization-aware knowledge distillation to transfer knowledge from the original model (i.e. teacher model) to the quantized one (i.e. student model). In the student model, the element in matrix $\mathbf{W}$ and vectors $\mathbf{g}$/$\mathbf{h}$ in Eq. (3) will be trained. We use cross-entropy based logits and mean-square-error based hidden state of the full-precision teacher model to direct the quantized student model [32]. Language modeling loss is not used. The cross-entropy is defined as
@@
where $\alpha$ is the hyper-parameter that balances the importance of the cross-entropy loss and the features in the intermediate layers. Please refer to A.6 for further discussions of this part.
@@ 4.1 Settings
For the training data of our quantization-aware knowledge distillation, we follow Liu et al. [21] to synthesize corpus using next token generation from the original teacher model. It randomizes the first token from vocabulary and generates the next token iteratively until reaching either the <EOS> token or the maximum length. Specially, the top-1 predictions are selected deterministically for the first 3 to 5 tokens, followed by stochastic sampling for the remaining tokens. We utilized LLaMA-7B to generate a total of 132k data entries, each with a maximum length of 2,048.
@@
Every KD experiment learns the training data over 50 epochs, from which 2048-token segments are selected. We employ NMF in scikit-learn  [^1] https://scikit-learn.org/ to decompose the weight matrices in SVID. The quantized student models are optimized by Adam [19] with $\beta_{1}=0.9$, $\beta_{2}=0.98$. The learning rate for all experiments is scheduled by cosine strategy. We use NVIDIA A100 GPUs and maintain FP16 precision while training quantized models. For additional details such as learning rate, please refer to Table 1.
Table 1: Training details of knowledge distillation.
Models learning rate $\alpha$ GPUs
@@
Wiki2 C4 Wino. Hella. PIQA BoolQ ARC-e ARC-c Avg.
@@
To our knowledge, there is no previous work exploring the 1-bit quantization of LLMs from a knowledge transfer perspective. To this end, we relax the quantization bit-width of baselines to 2 bits (W2A16) while maintaining the W1A16 setting in our method. We compare our method with GPTQ [14], LLM-QAT [21] and OmniQuant [30]. To ensure a fair comparison in terms of space usage, baselines do not employ grouped quantization. Additionally, we included the results of vanilla transformers with FP16 precision as a reference. While the recent work BitNet [38] also introduced one 1-bit model architecture, it only worked for training models from scratch. We also analyze its capability to transfer knowledge from the original models in Appendix A.5.
@@
Basically, we evaluate quantized models by testing the perplexity on the validation set, specifically on WikiText2 [23] and C4 [28]. Lower perplexity indicates that the compressed model is better at preserving the output distribution of the original model. Furthermore, accuracies of zero-shot tasks including Winograde [29], HellaSwag [41], PIQA [4], BoolQ [7], and ARC [8] are also reported. They evaluate if the capabilities of the original model on downstream tasks are retained. We utilize the open-sourced toolkit “LM-Evaluation-Harness” [^2] https://github.com/EleutherAI/lm-evaluation-harness to perform the perplexity test and all zero-shot tasks.